A 20 pC charge is placed at the origin and a -30 pC is places 3m to the right of it. What is the magnitude and direction of the electric field 7m meters from the origin?


1
Expert's answer
2021-02-28T07:37:22-0500

Answer

Electric field due to any point charge is given

"E=\frac{kq}{r^2}"

Electric field due 20 pC charge

"E_1=\frac{9\times10^9\times20\times10^{-12}}{7\times7}(-i)"


"=3.7\times10^{-3}(-i)" V/m


"E_2=\frac{9\times10^9\times30\times10^{-12}}{4\times4}(i)"

= "=16.9\times10^{-3}(i)" V/m

So resultant

The magnitude and direction of the electric field 7m meters from the origin

Is

E="E_1+E_2"

"=13.2\times10^{-3}(i)" v/m

Direction in positive x direction.